OpenStudy (anonymous):

A stone is tossed into a pound, creating a circular ripple. The radius of the ripple increases by 12cm/s. Write an equation that describes the ripple exactly 3s after the stone lands in the water.

OpenStudy (anonymous):

is this 10 grade math

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well i hope that helped

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

a) After 3 seconds, you have a radius of 3 * 12cm = 36 cm with its centre at the point where the stone entered the water. For a circle (x-h)^2 + (y-k)^2 = r^2, where (h,k) is the centre of the circle. Here, we've been told to call that point the origin (0,0), so the equation is x^2 + y^2 = (36)^2